In a copper wire of diameter 2mm, 3×10^23 electrons flow in one second. Calculate the electric field magnitude






1
Expert's answer
2022-03-02T14:39:35-0500

The electric field can be found from the Ohm's law in differential form:


"E = \dfrac{j}{\sigma}"

where "j" is the current density, and "\sigma=5.87\times 10^7S/m" is the conductivity of the copper. The current density is the following:



"j = \dfrac{I}{\pi d^2/4}"

where "d = 2\times 10^{-3}m" is the diameter and "\pi d^2/4" is the cross-sectional area. Thus, obtain:



"j = \dfrac{4\cdot 4.8\times 10^4}{\pi \cdot (2\times 10^{-3})} \approx 153A/m^2"


Thus, obtain:


"E = \dfrac{153A/m^2}{5.87\times 10^7S/m} \approx 2.6\times10^{-6}V/m"

Answer. "2.6\times10^{-6}V/m"